# and Reggio Emilia, Italy

### v2

@@ 7. Conclusion
We presented a comprehensive experimental comparison of the most used dataframe libraries to support practitioners in the selection of the most suitable solution to carry out their data preparation tasks on a single machine. To guarantee comparable results and increase usability, we developed Bento, a general framework for assessing the performance of dataframe libraries on four major data preparation stages: input/output (I/O), exploratory data analysis (EDA), data transformation (DT), and data cleaning (DC). We exploited Bento to perform a thorough comparative analysis of the libraries on heterogeneous datasets (varying in size, complexity, and features) previously adopted in literature and publicly available on Kaggle. In particular, we relied on the most voted Kaggle notebooks to assess the performance of the libraries on real-world data preparations pipelines validated by a large community. Further, we also compare their performance on the popular TPC-H benchmark to support our findings.
The results of our experiments show that there is no clear winner among the libraries, whose performance is strongly dependent on the size and features of the dataset at hand, on the specifications of the underlying machine, and on the kind of operations to carry out. Thus, we provided the readers with useful takeaways to support them in the selection of the right library for the data preparation task at hand. As future work, we plan to compare the libraries in a distributed environment and study the possible adoption of machine learning to suggest optimal library combinations based on datasets, tasks, hardware, and previous executions.
@@ Acknowledgements
We extend our appreciation to the master’s students of the 2021/2022 Big Data Management and Governance course at the University of Modena and Reggio Emilia for their contributions in testing the libraries. We warmly thank Leonardo S.p.A., co-funding the PhD program of Angelo Mozzillo. Further, this work was partially supported by the “Enzo Ferrari” Engineering Department (University of Modena and Reggio Emilia) within the projects FARD-2022 and FARD-2023, and by MUR within the project “Discount Quality for Responsible Data Science: Human-in-the-Loop for Quality Data” (code 202248FWFS).

### v3

@@ 7. Conclusion
We presented a comprehensive experimental comparison of the most used dataframe libraries to support practitioners in the selection of the most suitable solution to carry out their data preparation tasks on a single machine. To guarantee comparable results and increase usability, we developed Bento, a general framework for assessing the performance of dataframe libraries on four major data preparation stages: input/output (I/O), exploratory data analysis (EDA), data transformation (DT), and data cleaning (DC).
We exploited Bento to perform a thorough comparative analysis of the libraries on four heterogeneous datasets (varying in size, complexity, and features) previously adopted in literature and publicly available on Kaggle. In particular, we relied on three of the most voted Kaggle notebooks per dataset to assess the performance of the libraries on real-world data preparation pipelines validated by a large community. Further, we also evaluate their performance on the popular TPC-H benchmark to support our findings, including DuckDB to assess performance in comparison with an SQL-based system.
Our experimental evaluation shows that one size does not fit all when it comes to dataframe libraries for data preparation. However, we have distilled key takeaways that data practitioners can use as starting points for selecting the appropriate library for their task at hand.
As future work, we plan to compare the libraries in a distributed environment and study the possible adoption of machine learning to suggest optimal library combinations based on datasets, tasks, hardware, and previous executions. Further, we plan to explore how the libraries perform in combination with tools that optimize and rewrite entire dataframe-based data preparation pipelines, such as Dataprep (Peng et al., 2021) and Dias (Baziotis et al., 2024).
@@ Acknowledgements
We extend our appreciation to the master’s students of the 2021-2022 Big Data Management and Governance course at the University of Modena and Reggio Emilia for their contributions in testing the libraries. We warmly thank Leonardo S.p.A., co-funding the PhD program of Angelo Mozzillo. Further, this work was partially supported by the “Enzo Ferrari” Engineering Department (University of Modena and Reggio Emilia) within the projects FARD-2022 and FARD-2023, and by MUR within the project “Discount Quality for Responsible Data Science: Human-in-the-Loop for Quality Data” (code 202248FWFS).